# Global magnetohydrodynamic simulations of the inner regions of protoplanetary discs. I. Zero-net flux regime

### v1

@@ 5.2.3 Long-term evolution
Following the merger process, the weakly coherent $m=1$ vortex shears out over a timescale of $\sim 40t_{\text{DZI}}$, with its power budget redistributed into the axisymmetric ($m=0$) component. This timescale is comparable to that reported by Flock et al. (2015), who used the same azimuthal domain extent and in their study of the outer dead zone edge.
@@
To evaluate the influence of numerical resolution on the rapidity of this decay, we performed a resolution test (ZNF-HRES), with double the number of cells in azimuth and radius. We then analyse the Fourier power spectrum of the density field, focusing on the $m=1$ mode, $P_{\rho,m=1}$, defined in Section 3.4.2, and present the results in Fig. 18.
This reveals an order-of-magnitude difference in the evolution of the $m=1$ power across resolutions, highlighting the sensitivity of vortex persistence to numerical resolution at this level.
@@
We also consider how vortex evolution may proceed in the high-resolution limit of global simulations. Multiple vortex stages observed in our simulations are likely susceptible to different variants of the elliptical instability (Pierrehumbert, 1986; Kerswell, 2002; Lesur & Papaloizou, 2009; Mizerski & Bajer, 2009; Railton & Papaloizou, 2014): the three-dimensional $m=7$ vortices exhibit relatively low aspect ratios ($\sim 3$), whereas the later stage $m=1$ vortex displays higher aspect ratios ($>\!5$). Whilst we see no evidence of elliptical instability at the resolutions or timescales considered, it may emerge at higher resolutions or over longer integration times (e.g. Richard et al., 2013). If triggered, it could render the vortex cores turbulent, disrupting dust aggregation at their centres and potentially destroying the coherent vortical streamlines.
@@
Figure 18: Fourier power time series of the $m=1$ density mode, $P_{\rho,m=1}\,(R=1.05,\text{au},t)$, evaluated at the vortex centres. The power (and $m=1$ vortex) decays more rapidly in ZNF-FID (solid), compared to the higher resolution test ZNF-HRES (dotted).
@@
To investigate the potential for vortex regeneration, we extend the runtime in ZNF-LONG to $225t_{\text{DZI}}$. Whilst renewed vortensity activity appeared at the dead–active zone interface, following the dissipation of the initial vortex, no subsequent development of a vortensity structure capable of triggering the RWI occurred. A likely explanation is that the relatively small active zone allows the inner boundary condition to interfere with the long-term evolution and radial redistribution of vortensity, both essential for re-establishing the conditions necessary to trigger the RWI. Radially extended simulations, such as the vertical-net flux models presented in the second paper of this series and the outer dead zone study by Flock et al. (2015), show evidence for non-migrating periodic vortex life cycles.
@@
Figure 19: Magnetic field evolution for ZNF-FID, averaged in azimuth and over a single orbit at the dead–active zone interface, as indicated above each plot. The background colour shows the (normalised) toroidal magnetic field $R\langle B_{\phi}\rangle_{\phi,t}$ and the dashed white lines mark the disc–corona transition. The poloidal magnetic field $\langle\mathbf{B_{p}}\rangle_{\phi,t}$ morphology, shown using black field lines and a LIC overlay in greyscale, reveal coherent large-scale clockwise loops in the dead zone and MRI active region (see Fig. 5). Note the broad anti-correlation between the sign of the toroidal and radial magnetic fields, and that the disordered poloidal field in the dead zone, in the left panel, is extremely weak.
We complete the presentation of our results by outlining the evolution of magnetic field structures in the dead zone in Section 6.1. In Section 6.2 we examine the coherent poloidal magnetic field structures in the dead zone, and their capacity to drive a weak magnetic pressure wind. Finally, in Section 6.3 we speculate on the formation mechanism of these complex magnetic field structures in the dead zone.
@@ 6.1 Overview
We begin by briefly examining the global evolution of the magnetic field in the dead zone in ZNF-FID. An initially toroidal magnetic field confined to the active zone (see left panel of Fig. 3), expands and evolves into a complex, large-scale field structure with both toroidal and poloidal components, which gradually fills the entire extent of the dead zone.
To capture this evolution in more detail, Fig. 19 presents three snapshots of the field evolution, averaged azimuthally and temporally over a single orbit at the dead–active zone interface. Whilst Fig. 3 only illustrates the spreading of the toroidal field, the addition of poloidal field lines $\langle\mathbf{B_{p}}\rangle_{\phi,t}$ and their associated LIC overlay, in Fig. 19 provides a complementary perspective, highlighting the emergence and organisation of the poloidal component. The apparently disordered poloidal field present in the dead zone at early times (left panel of Fig. 19), is extremely weak and is the result of the visualisation technique.
Notably, the growth of $B_{\phi}$ in the dead zone coincides with the emergence of increasingly ordered poloidal field structures. In particular, we observe a tendency toward anti-correlation between the signs of $B_{r}$ and $B_{\phi}$, along with the formation of concentric, coherent large-scale poloidal loops that gradually diffuse outward over time and encircle an inversion point (see middle panel of Fig. 19). These features are examined in greater detail in the following two sections.
@@ 6.2.1 Morphology and evolution
The initially absent poloidal magnetic field in the dead zone gradually develops into a striking and complex configuration by $t_{\mathrm{in}}\sim 250$, as shown in the middle panel of Fig. 19. This consists of large-scale clockwise poloidal loops encircling an inversion point, with ordered vertical field oriented positively on the interior and negatively on the exterior. Throughout the rest of the simulation, this configuration migrates outward, and becomes increasingly coherent: by $t_{\text{in}}\sim 730$ the large-scale clockwise poloidal field loops are clearly identifiable in the dead zone.
@@
In Fig. 20 we present $(R,t)$ spacetime plots to characterise the radial structure of the vertical magnetic field in the disc. The top panel shows $R\langle{B_{z}}\rangle_{\phi,1H}$, the vertical field, averaged in azimuth and the meridional direction over a single scale height $H$ either side of the midplane. This confirms the emergence of an ordered vertical field geometry that extends around the midplane: forming coherent loops in both the active zone (see Fig. 5) and the dead zone, where the loops retain the same orientation, as discussed above.
@@
This ordered vertical field in the dead zone migrates outward, with the inversion point developing near the dead–active zone interface, as evident in both panels of Fig. 20. The bottom panel shows $\beta_{z}^{\text{coh}}(z=0)$, the coherent vertical component of the plasma-$\beta$ parameter defined in equation (25). The coherent field structure, which we characterise by the inversion point, is transported radially outward at a velocity of $\sim 5\times 10^{-4}u_{K}$ at the midplane, based on a visual match of its trajectory, in code units where $R_{0}=1$. This value is broadly consistent with a characteristic Ohmic diffusion speed, $v_{O}\sim\eta_{O}/R$, yielding $v_{O}\sim(H/R)^{2}\,\text{R}_{m}^{-1}\,u_{K}\sim 10^{-4}u_{K}$. Also see top-right panel of fig. 11 in Lesur (2021b) and note that their velocities are normalised by an extra factor of $\varepsilon^{-1}$. This suggests that the midplane outward migration is driven by non-ideal MHD diffusion, dominated by the Ohmic component.
@@
Figure 20: Spacetime $(R,t)$ diagrams of the vertical magnetic field evolution. Top: Azimuthal azimuthal and meridional integration of (normalised) vertical magnetic field, $R\langle B_{z}\rangle_{\phi,1H}$. The trend of upward oriented field (orange) in the inner active/dead zone and downward oriented field (blue) in the outer active/dead zone, is consistent with the large-scale poloidal field loops of the same orientation in the MRI active (Fig. 5) and MRI dead (Fig. 19) zones. Bottom: The coherent vertical component of the plasma-$\beta$ parameter, $\beta_{z}^{\mathrm{coh}}$, at the disc midplane reveals the outward migration of field structures from the dead–active zone interface (dashed white line). The inversion point is well matched by an outward propagation speed of $5\times 10^{-4}u_{K}$ (dot-dashed black line), based on a visual fit.
@@ 6.2.2 Outflow characteristics in the inner dead zone
As this complex poloidal field structure expands outward via non-ideal MHD diffusion, the region of the MRI dead disc threaded by a coherent, vertically oriented field increases (see top panel of Fig. 20). By $t_{\mathrm{in}}\sim 700$, the coherent, positively oriented vertical magnetic field extends across the full vertical extent of the disc and spans the radial range, $1\,\mathrm{au}\lesssim r\lesssim 1.4\,\mathrm{au}$, as shown in the right panel of Fig. 19. The magnetic field topology in this region closely resembles that of vertical-net flux global non-ideal MHD simulations (e.g. fig. 2 in Gressel et al. 2020), which impose a strong vertical laminar torque on the disc, sufficient to drive accretion through the dead zone. In the remainder of this section, we characterise the magnetic field structure along the spherical contour $r=1.15\,\text{au}$, and evaluate its ability to impose vertical torques and drive accretion, in analogy with vertical-net flux configurations.
@@
The relative strength of the vertical magnetic field is quantified using the coherent vertical component of the plasma-$\beta$ parameter, defined in equation (25). At $r=1.15\,\text{au}$, $\beta_{z}^{\text{coh}}\sim 5\times 10^{5}$ for the configuration shown in the right panel of Fig. 19, indicating a weak but dynamically relevant vertical field. Indeed, the axisymmetric non-ideal simulations of Lesur (2021b) conclude that such $\beta_{z}^{\text{coh}}$ values can drive weak magnetic pressure winds.
@@
To assess whether this magnetic field configuration can drive a magnetic outflow, we first examine the poloidal velocity $\langle\mathbf{u_{p}}\rangle_{\phi,t}$ and magnetic energy $\langle B^{2}\rangle_{\phi,t}$, over the spherical contour $r=1.15\,\mathrm{au}$, averaged in azimuth and over the interval $t_{\mathrm{in}}\in[600,700]$. As shown in Fig. 21, the flow is directed vertically away from the midplane in the regions where $|z|\gtrsim 1H$. Just below the disc’s surface layers, the poloidal velocity is $u_{p}\sim 2\times{10}^{-2}$, equivalent to the local sound speed in the disc, $c_{s}=\varepsilon R^{-1/2}$. Meanwhile, Fig. 22 shows that the toroidal component of the magnetic energy $\langle B_{\phi}^{2}\rangle_{\phi,t}$ dominates over the other components, reaching a maximum on both sides of the disc midplane, while decreasing sharply in the regions $|z|\gtrsim 2.5H$. This configuration supports vertical acceleration by laminar magnetic pressure and marks the regions around $|z|\approx 2.5H$ where the weak magnetic field extracts and stores angular momentum. The vertical asymmetry in the velocity and magnetic field structure is driven by the diffuse current sheet adopting an off-midplane configuration (as shown in the right panel of Fig. 19).
@@
Furthermore, the vertical flow in the inner dead zone is accompanied by a comparatively weak midplane inward radial flow on the order of $\sim 0.01c_{s}$, consistent with the midplane plots presented in Fig. 10. To assess the magnetic contribution to the angular momentum extraction driving this accretion flow, we quantify the coherent magnetic stresses, defined in equation (23), along the contour $r=1.15$ au. At $z=\pm 3H$, the coherent components of the (normalised) Maxwell stress are, $\mathcal{M}^{\text{coh}}_{\theta\phi}/\langle P\rangle_{\phi,t}\sim\mathcal{M}^{\text{coh}}_{r\phi}/\langle P\rangle_{\phi,t}\sim 2\times 10^{-3}$. An additional source of radial angular momentum transport arises from correlated velocity fluctuations near the disc midplane, yielding $\mathcal{R}_{r\phi}/\langle P\rangle_{\phi,t}\sim 5\times 10^{-4}$, as discussed in Section 4.4. The contribution of each stress component to the mass accretion rate is discussed in Lesur (2021a), but is unable to be applied exactly to this evolving, radially restricted region. Nonetheless, within the thin-disc approximation, vertical stresses contribute $\sim\!R/H$ more to the accretion rate than the vertically integrated radial stresses. Therefore, coherent vertical angular momentum extraction is a non-negligible driver of the accretion at this radius.
@@
In summary, we attribute the vertical acceleration of the flow in this region to a weak magnetothermal wind (e.g. Bai et al., 2016); driven by a combination of the vertical laminar magnetic pressure gradient, $\partial_{z}(B^{2})/8\pi$, which is dominated by $B_{\phi}$, and the thermal pressure acceleration provided by the hot corona, closer to the disc surfaces.
@@
Figure 21: Decomposition of the poloidal velocity field $\langle\mathbf{u_{p}}\rangle_{\phi,t}$ into its coordinate components along the spherical contour $r=1.15\,\text{au}$. This contour corresponds with the coherent vertical magnetic field configuration at the inner edge of the dead zone presented in Fig. 19, and the components have been averaged in azimuth and over the interval $t_{\mathrm{in}}\in[600,700]$. The vertical flow (dotted orange line) is directed away from the midplane, which is driven by the magnetic pressure configuration presented in Fig. 22.
@@
Figure 22: Decomposition of the magnetic energy $\langle B^{2}\rangle_{\phi,t}$ into its coordinate components along the spherical contour $r=1.15\,\text{au}$. This contour corresponds with the coherent vertical field configuration at the inner edge of the dead zone presented in Fig. 19, and the components have been averaged in azimuth and over the interval $t_{\mathrm{in}}\in[600,700]$. The grey shaded regions indicate where the vertical magnetic pressure gradient (dominated by $B_{\phi}^{2}$), acts to accelerate the flow away from the disc.

### v3

@@ 5.2.3 Long-term evolution
Following the merger process, the weakly coherent $m=1$ vortex shears out over a timescale of $\sim\!40t_{\text{DZI}}$, with its power budget redistributed into the axisymmetric ($m=0$) mode. This timescale is comparable to that reported by flock_gaps_2015, who used the same azimuthal domain extent in their outer dead zone edge study.
@@
To evaluate the influence of numerical resolution on the rapidity of this decay, we performed a resolution test (ZNF-HRES), with double the number of cells in azimuth and radius. We then analyse the Fourier-power spectrum of the density field, focusing on the $m=1$ mode, $P_{\rho,m=1}$, defined in Section 3.4.2, and present the results in Fig. 18. This reveals an order-of-magnitude difference in the evolution of the $m=1$ power across resolutions, highlighting the sensitivity of vortex persistence to numerical resolution at this level.
@@
We also consider how vortex evolution may proceed in the high-resolution limit of global simulations. Multiple vortex stages observed in our simulations are likely susceptible to different variants of the elliptical instability (pierrehumbert_universal_1986; kerswell_elliptical_2002; lesur_stability_2009; mizerski_magnetoelliptic_2009; railton_local_2014): the three-dimensional $m=7$ vortices exhibit relatively low aspect ratios ($\sim\!3$), whereas the later-stage $m=1$ vortex displays higher aspect ratios (>5>\!5). Whilst we see no evidence of elliptical instability at the resolutions or timescales considered, it may emerge at higher resolutions or over longer integration times (e.g. richard_structure_2013). If triggered, it could render the vortex cores turbulent, disrupting dust aggregation at their centres and potentially destroying the coherent vortical streamlines.
@@
Figure 18: Fourier-power time series of the $m=1$ density mode, $P_{\rho,m=1}\,(R=1.05,\text{au},t)$, evaluated at the vortex centres. The power (and $m=1$ vortex) decays more rapidly in ZNF-FID (solid), compared to the higher-resolution test ZNF-HRES (dotted).
@@
To investigate the potential for vortex regeneration, we extend the runtime in ZNF-LONG to $225\,t_{\text{DZI}}$. Whilst renewed vortensity activity appeared at the dead–active zone interface, following the dissipation of the initial vortex, no subsequent development of a vortensity structure capable of triggering the RWI occurred. A likely explanation is that the relatively small active zone allows the inner boundary condition to interfere with the long-term evolution and radial redistribution of vortensity, both essential for re-establishing the conditions necessary to trigger the RWI. Radially extended simulations, such as the outer dead zone study by flock_gaps_2015, show evidence for non-migrating periodic vortex life cycles.
Figure 19: Magnetic field evolution for ZNF-FID, averaged in azimuth and over a single orbit at the dead–active zone interface, as indicated above each plot. The background colour shows the normalised toroidal magnetic field, $R\langle B_{\phi}\rangle_{\phi,t}$, and the dashed white lines mark the disc–corona transition. The poloidal magnetic field, $\langle\mathbf{B_{p}}\rangle_{\phi,t}$, shown using black field lines and a LIC overlay in greyscale, reveals coherent large-scale clockwise loops in the dead zone and MRI-active region (see Fig. 5). Note the broad anti-correlation between the sign of the toroidal and radial magnetic fields, and that the disordered poloidal field in the dead zone in the left panel is extremely weak.
@@
We complete the presentation of our results by outlining the evolution of magnetic field structures in the dead zone in Section 6.1. In Section 6.2 we examine the coherent poloidal magnetic field structures in the dead zone, and their capacity to drive a weak magnetic-pressure wind. Finally, in Section 6.3 we speculate on the formation mechanism of these magnetic field structures in the dead zone.
@@ 6.1 Overview
We begin by briefly examining the global evolution of the magnetic field in the dead zone in ZNF-FID. An initially toroidal magnetic field confined to the active zone (see left panel of Fig. 3), expands and evolves into a complex, large-scale field structure with both toroidal and poloidal components, which gradually fills the entire dead zone.
To capture this evolution in more detail, Fig. 19 presents three snapshots of the field evolution, averaged azimuthally and temporally over a single orbit at the dead–active zone interface. Whilst Fig. 3 only illustrates the spreading of the toroidal field, the addition of poloidal magnetic field lines, $\langle\mathbf{B_{p}}\rangle_{\phi,t}$, and their associated LIC overlay, in Fig. 19 provides a complementary perspective, highlighting the emergence and organisation of the poloidal component. Please note that the apparently disordered poloidal field present in the dead zone at early times (left panel of Fig. 19), is extremely weak and an unfortunate result of the visualisation technique.
Notably, the growth of $B_{\phi}$ in the dead zone coincides with the emergence of increasingly ordered poloidal field structures. In particular, we observe a tendency towards anti-correlation between the signs of $B_{r}$ and $B_{\phi}$, along with the formation of concentric, coherent large-scale poloidal loops that gradually diffuse outward over time and encircle an inversion point (see middle panel of Fig. 19). These features are examined in greater detail in the following two sections.
@@ 6.2.1 Morphology and evolution
The initially absent poloidal magnetic field in the dead zone gradually develops into a striking and complex configuration by $t_{\mathrm{in}}\!\sim\!250$, as shown in the middle panel of Fig. 19. This consists of large-scale, clockwise poloidal loops encircling an inversion point, with ordered vertical field oriented positively on the interior and negatively on the exterior. Throughout the rest of the simulation, this configuration migrates outward, and becomes increasingly coherent: by $t_{\text{in}}\!\sim\!730$, the large-scale clockwise poloidal field loops are clearly identifiable in the dead zone.
@@
In Fig. 20, $(R,t)$ spacetime plots are presented to characterise the radial structure of the vertical magnetic field in the disc. The top panel shows $R\langle{B_{z}}\rangle_{\phi,1H}$, the vertical field averaged in azimuth and over the meridional direction within a single scale height, $H$, either side of the midplane. This confirms the emergence of an ordered vertical field geometry that extends around the midplane, forming coherent loops in both the active zone (see Fig. 5) and the dead zone, where the loops retain the same orientation, as discussed above.
@@
This ordered vertical field in the dead zone migrates outward, with the inversion point developing near the dead–active zone interface, as evident in both panels of Fig. 20. The bottom panel shows $\beta_{z}^{\text{coh}}(z=0)$, the coherent vertical component of the plasma-$\beta$ parameter, as defined in equation (24). The coherent field structure, which we characterise by the inversion point, is transported radially outward at a velocity of roughly $5\times 10^{-4}u_{\text{K}}$ at the midplane, based on a visual match of its trajectory. This value is broadly consistent with a characteristic Ohmic-diffusion speed, $v_{O}\sim\eta_{\text{O}}/R$, yielding $v_{\text{O}}\!\sim\!(H/R)^{2}\,\text{R}_{\text{m}}^{-1}\,u_{\text{K}}\!\sim\!10^{-4}u_{\text{K}}$. Also see top-right panel of fig. 11 in lesur_systematic_2021 and note that their velocities are normalised by an extra factor of $\varepsilon^{-1}$. This suggests that the midplane outward migration is driven by non-ideal MHD diffusion, dominated by the Ohmic component.
@@
Figure 20: Spacetime $(R,t)$ diagrams of the vertical magnetic field evolution. Top: Azimuthally and meridionally integrated (normalised) vertical magnetic field, $R\langle B_{z}\rangle_{\phi,1H}$. The trend of upward-oriented field (orange) in the inner active/dead zone and downward-oriented field (blue) in the outer active/dead zone is consistent with the large-scale poloidal field loops of the same orientation in the MRI-active (see Fig. 5) and MRI-dead (see Fig. 19) zones. Bottom: The coherent vertical component of the plasma-$\beta$ parameter, $\beta_{z}^{\mathrm{coh}}$, at the midplane reveals the outward migration of field structures from the dead–active zone interface (dashed white line). The inversion point is well matched by an outward propagation speed of $5\times 10^{-4}u_{\text{K}}$ (dot-dashed black line), based on a visual fit.
@@ 6.2.2 Outflow characteristics in the inner dead zone
As this complex poloidal field structure expands outward via non-ideal MHD diffusion, the region of the MRI-dead disc threaded by a coherent, vertically oriented field increases (see top panel of Fig. 20). By $t_{\mathrm{in}}\!\sim\!700$, the coherent, positively oriented vertical magnetic field extends across the full vertical extent of the disc and spans the radial range, $1\,\mathrm{au}\lesssim r\lesssim 1.4\,\mathrm{au}$, as shown in the right panel of Fig. 19. The magnetic field topology in this region closely resembles that of vertical-net flux, global, non-ideal MHD simulations (e.g. fig. 2 in gressel_global_2020), which impose a strong vertical laminar torque on the disc, sufficient to drive accretion through the dead zone. In the remainder of this section, we characterise the magnetic field structure along the spherical contour $r=1.15\,\text{au}$, and evaluate its ability to impose vertical torques and drive accretion, in analogy with vertical-net flux configurations.
@@
The relative strength of the vertical magnetic field is quantified using the coherent vertical component of the plasma-$\beta$ parameter, defined in equation (24). At $r=1.15\,\text{au}$, $\beta_{z}^{\text{coh}}\!\sim\!5\times 10^{5}$ for the configuration shown in the right panel of Fig. 19, indicating a weak but dynamically relevant vertical field. Indeed, the axisymmetric non-ideal MHD simulations of lesur_systematic_2021 conclude that such $\beta_{z}^{\text{coh}}$ values can drive weak magnetic-pressure winds.
@@
To assess whether this magnetic field configuration drives a magnetic outflow, we first examine the poloidal velocity, $\langle\mathbf{u_{p}}\rangle_{\phi,t}$, and magnetic energy, $\langle B^{2}\rangle_{\phi,t}$, over the spherical contour $r=1.15\,\mathrm{au}$, averaged in azimuth and over the interval $t_{\mathrm{in}}\in[600,700]$. As shown in Fig. 21, the flow is directed vertically away from the midplane in the regions where $|z|\gtrsim 1H$. Just below the disc’s surface layers, the poloidal velocity is $u_{\mathrm{p}}\!\sim\!2\times{10}^{-2}$, equivalent to the local sound speed in the disc, $c_{\text{s}}=\varepsilon R^{-1/2}$. Meanwhile, Fig. 22 shows that the toroidal component of the magnetic energy, $\langle B_{\phi}^{2}\rangle_{\phi,t}$, dominates over the other components, reaching a maximum on both sides of the disc midplane, whilst decreasing sharply in the regions $|z|\gtrsim 2.5H$. This configuration supports vertical acceleration by laminar magnetic pressure and identifies the regions around $|z|\approx 2.5H$ as where the weak magnetic field extracts and stores angular momentum. Note that the diffuse off-midplane current sheet associated with this structure (shown in the right panel of Fig. 19) imprints a vertical asymmetry on the velocity and magnetic field profiles in Fig. 21.
@@
Furthermore, the vertical flow in the inner dead zone is accompanied by a comparatively weak midplane inward radial flow on the order of $\sim\!0.01c_{\text{s}}$, consistent with the midplane plots presented in Fig. 10. To assess the magnetic contribution to the angular-momentum extraction driving this accretion flow, we quantify the coherent magnetic stresses, defined in equation (22), along the contour $r=1.15$ au. At $z=\pm 3H$, the coherent components of the (normalised) Maxwell stress are, $\mathcal{M}^{\text{coh}}_{\theta\phi}/\langle P\rangle_{\phi,t}\!\sim\!\mathcal{M}^{\text{coh}}_{r\phi}/\langle P\rangle_{\phi,t}\!\sim\!2\times 10^{-3}$. An additional source of radial angular momentum transport arises from correlated velocity fluctuations near the disc midplane, yielding $\mathcal{R}_{r\phi}/\langle P\rangle_{\phi,t}\!\sim\!5\times 10^{-4}$, as discussed in Section 4.4. The contribution of each stress component to the mass accretion rate is discussed in lesur_magnetohydrodynamics_2021, but is unable to be applied exactly to this evolving, radially restricted region. Nonetheless, within the thin-disc approximation, vertical stresses contribute $\sim\!R/H$ more to the accretion rate than vertically integrated radial stresses. Therefore, coherent vertical angular-momentum extraction is a non-negligible driver of accretion at this radius.
@@
In summary, vertical acceleration of the flow in this region is due to a weak magnetothermal wind (e.g. bai_magneto-thermal_2016); driven by a combination of the vertical laminar magnetic pressure gradient, $\partial_{z}(B^{2})/8\pi$, which is dominated by $B_{\phi}$, and the thermal pressure acceleration provided by the hot corona, closer to the disc surfaces.
@@
Figure 21: Decomposition of the poloidal velocity field, $\langle\mathbf{u_{p}}\rangle_{\phi,t}$, into its coordinate components along the spherical contour $r=1.15\,\text{au}$. This contour corresponds with the coherent vertical magnetic field configuration at the inner edge of the dead zone presented in Fig. 19, and the components have been averaged in azimuth and over the interval $t_{\mathrm{in}}\in[600,700]$. The vertical flow (dotted orange line) is directed away from the midplane, which is driven by the magnetic pressure configuration presented in Fig. 22.
@@
Figure 22: Decomposition of the magnetic energy, $\langle B^{2}\rangle_{\phi,t}$, into its coordinate components along the spherical contour $r=1.15\,\text{au}$. This contour corresponds with the coherent vertical field configuration at the inner edge of the dead zone presented in Fig. 19, and the components have been averaged in azimuth and over the interval $t_{\mathrm{in}}\in[600,700]$. The grey-shaded regions indicate where the vertical magnetic pressure gradient (dominated by $B_{\phi}^{2}$) accelerates the flow away from the disc.